I am stuck on finding a quadratic model for (1,-2), (2,-2), (3,-4). I do not know how to put x numbers and y numbers in the formulas and how to solve the answer? How do you get the three equations with variables a, b, and c?
How do you find the values for a,b, and c?
Thanks!

Answer by scott8148(6628)

general equation ___ y = ax^2 + bx + c

in this case:

for (1,-2) ___ -2 = a(1)^2 + b(1) + c ___ a + b + c = -2

for (2,-2) ___ -2 = a(2)^2 + b(2) + c ___ 4a + 2b + c = -2

for (3,-4) ___ -4 = a(3)^2 + b(3) + c ___ 9a + 3b + c = -4

three equations and three unknowns